• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gilman

    Log entry time 15:59:54 on November19,2008

    Entry number 249753

    keyword=day shift summary - 19 Nov 2008



    Left spectrometer vacuum gauge increases from 8 to 9 as we start shift.

    Run 3662 in progress at start of shift.

    run start stop events comments
    3662 ---- 8:49 1 M
    3663 8:50 9:10 0.3 M stop run to reset hrsl vacuum pump

    9:10 ask for controlled access
    9:41 access over, ask for beam back
    9:52 MCC calls and says they are sending beam back

    3664 9:51 10:41 0.2 M beam mostly off during this run
    10:01: ioc issue, expect 10-15 minutes off...
    11:43: beam returns!

    3665 11:43 12:54 1 M ok, beam ratty, many long trips
    3666 12:55 13:43 1 M ok, ~11 beam trips in last hour
    3667 13:44 14:38 1 M ok. ~18 beam trips
    3668 14:38 15:25 1 M ok. ~10 trips, one 3-4 minutes long
    3669 15:26 ~15:50 435 k stopped when Q1 left trips.

    Magnet still off at end of shift.